OpenStudy (anonymous):

sina+sinb=sqrt of 3/2,cosa+cosb=sqrt of 1/2 then 'a+b'?

OpenStudy (raden):

use the identity : cos(a+b) = cosacosb - sinasinb let's see what happend if you subtitute all the values given

OpenStudy (anonymous):

so far i got this sinasinb+cosacosb=0

OpenStudy (raden):

that's not i mean first, distribute all values given into formula above

OpenStudy (raden):

opsss, i read the equation wrong. sorry, may bad

OpenStudy (anonymous):

I have no idea how to approach this question, I thought of squaring the two equations and adding them together so I can cancel out some values and I forgot to mention, value of a and b is larger than 0 and smaller than 180

OpenStudy (raden):

yes, that's nice idea! sina+sinb=sqrt of 3/2 -> sin^2 a + sin^2 + 2sinasinb = 3/2 or =>1 + 2sinasinb = 3/2 => 2sinasinb = 3/2 - 1 => 2sinasinb = 1/2 => sinasinb = 1/4 cosa+cosb=sqrt of 1/2 --> cos^2 a + cos^2 b + 2cosacosb = 1/2 or =>1 + 2cosacosb = 1/2 => 2cosacosb = 1/2 - 1 => 2cosacosb = -1/2 => cosacosb = -1/4 then use the identity like i said above : cos(a+b) = cosacosb - sinasinb therefore, cos(a+b) = -1/4 - 1/4 cos(a+b) = -1/2 cos(a+b) = cos120 so, a+b = 120 with a,b < 180

OpenStudy (raden):

i hope that make sense :)

OpenStudy (anonymous):

hm I understand everything but Im not sure where this 1 came from.. 1 + 2sinasinb = 3/2

OpenStudy (anonymous):

Hint: Use this identity then divide equations. sinA + sinB = sin(A+B/2)cos(A-B/2) cosA + cosB = cos(A+B/2)cos(A-B/2)

OpenStudy (raden):

hmmm.. yes, that still wrong. i forgot to add both equation correction ; sin^2 a + sin^2 b + 2sinasinb + cos^2 a + cos^2 b + 2cosacosb = 3/2 + 1/2 (sin^2 a + cos^2 a) + (sin^2 b + cos^2 b) + 2(sinasinb + cosacosb) = 2 1+1 + 2cos(a-b) = 2 2 + 2cos(a-b) = 2 2cos(a-b) = 2 - 2 2cos(a-b) = 0 cos(a-b) = 0 cos(a-b) = cos90 a-b = 90 hmmm... where a+b ? --"

OpenStudy (anonymous):

@SAValkyrie although squaring is fine sometimes, it should be avoided as far as possible as it'll lead to surplus solutions. You'll have to plug back the values and crosscheck.

OpenStudy (anonymous):

so you would get tan(A+B/2)=sqrt(3)

OpenStudy (anonymous):

YEP. And you know when tanC = sqrt3 It means its in the first quad, And it is equal to 60 degrees. So A+B = 120 degrees

OpenStudy (anonymous):

Oh I see that answer makes sense, although I tried doing the squaring thingy because in the solution i got, it said to square both equations and add them together. But that was all that was written in the solution so I was wondering if there's a way to solve this without using different identities

OpenStudy (anonymous):

I am actually in Grade 10 so I havent learned those identities in school yet, the question was actually from math competition.. ;;

OpenStudy (anonymous):

Hm Thanks for all the help, I guess I would have to work on memorizing the identities of the trigonometry from now on!!

OpenStudy (raden):

ah i got it now we have a-b = 90 ------> a = b+90 take the 1st equ : sina+sinb=sqrt of 3/2 subtitute a = b + 90 sin(b+90) + sinb = sqrt of 3/2 cosb + sinb = sqrt(3/2) take the 2nd equ : cosa+cosb=sqrt of 1/2 cos(b+90) + cosb = sqrt(1/2) -sinb + cosb = sqrt(1/2) divide both equation, we get (cosb + sinb)/ (-sinb + cosb) = sqrt(3/2)/ sqrt(1/2) (cosb + sinb)/ (cosb - sinb) = sqrt(3) cosb + sinb = sqrt(3)(cosb - sinb) cosb + sinb = sqrt(3)cosb -sqrt(3) sinb (1+sqrt(3))sinb = (1+sqrt(3))cosb sinb/cosb = 1 tanb = 1 b = 45 a-b = 90 a - 45 = 90 a = 45 + 90 a = 135 so, a+b = 135 + 45 = 180

OpenStudy (anonymous):

\[\sin a+\sin b=\sqrt{\frac{ 3 }{ 2 }},\cos a+\cos b=\sqrt{\frac{ 1 }{ 2}}\] squaring and adding, \[\sin ^{2}a+\sin ^{2}b+2\sin a \sin b+\cos ^{2}a+\cos ^{2}b+2\cos a \cos b=\frac{ 3 }{2 }+\frac{ 1 }{2 }=2\] \[\left( \sin ^{2}a+\cos ^{2}a \right)+\left(\sin ^{2}b+\cos ^{2}b \right)+2\left( \cos a \cos b+\sin a \sin b \right)=2\] \[1+1+2\cos \left( a-b \right)=2\] \[\cos \left( a-b \right)=0\] Again \[\left( \cos ^{2}a-\sin ^{2}a \right)+\left( \cos ^{2}b-\sin ^{2} b\right)+2\left( \cos a \cos b-\sin a \sin b \right)=\frac{ 1 }{2 }-\frac{ 3 }{ 2}=-1\] \[\cos 2a+\cos 2b+2\cos \left( a+b \right)=-1\] \[2\cos \frac{ 2a+2b }{ 2 }\cos \frac{ 2a-2b }{2 }+2\cos \left( a+b \right)=-1\] \[2\cos \left( a+b \right)\cos \left( a-b \right)+2\cos \left( a+b \right)=-1\] 0+2cos(a+b)=-1 \[\cos \left( a+b \right)=\frac{ -1 }{2 }=-\cos \frac{ \pi }{ 3 }=\cos \left( \pi-\frac{ \pi }{ 3 } \right)=\cos \frac{ 2\pi }{ 3 }\] \[a+b=\frac{ 2\pi }{ 3}\]

OpenStudy (anonymous):

cosacosb+sinasinb how did this became cos(a-b)?? Is that one of the identities?

OpenStudy (anonymous):

cos(a-b)=cosacosb+sinasinb

OpenStudy (anonymous):

Oh I see, thanks for the answer, sadly I can only choose 1 best response D:

OpenStudy (raden):

i gave for surjithayer, and you have to give to @shiddantsharan :)